A transformer has the voltages 25 V_pri and 140 V_11 12 13 14
Answer #1

(11) The current flowing in the secondary which is given as :

using a formula, we have

IS = IP (VS / VP) = (3.5 mA) [(140 V) / (25 V)]

IS = 19.6 mA

(12) A machine in the factory needs 31 V to operate.

a. The voltage of the electrical power system in the factory will be given as :

using a formula, we have

VS = VP (NS / NP)

(31 V) = VP (0.0675)

VP = (31 V) / (0.0675)

VP = 459.2 V

(13) The inducatnce of the coil which will be given as :

using a formula, we have

L = \mu0 n2 L A

where, A = area of the coil = \pir2

L = (4\pi x 10-7 mT/A) (180)2 (0.88 m) [(3.14) (0.6 m)2]

L = 404808.3 x 10-7 H

L = 40.5 x 10-2 H

L = 40.5 mH

(14) Given that :

inductance, L = 0.018 H

number of turns, n = 3100

diameter of cylinder, D = 2.6 cm

radius of cylinder, r = 1.3 cm = 0.013 m

a. It's length will be given as :

L = \mu0 n2 L A

(0.018 H) = (4\pi x 10-7 mT/A) (3100)2 L [(3.14) (0.013 m)2]

(0.018 H) = (64051.5 x 10-7 m3 T/A) L

L = (0.018 H) / (64051.5 x 10-7 m3 T/A)

L = 2.81 m

convert m into cm :

L = 281 cm